OpenStudy (destinymasha):

Evaluate the following absolute value function for x = - 3. F [x] = 2 | 4x - 5| + 2

OpenStudy (mathstudent55):

First, substitute -3 for x. Then calculate the expression inside the absolute sign. Then take the absolute value. Then multiply by 2. Then add 2.

OpenStudy (destinymasha):

ok, hold on a sec

OpenStudy (mathstudent55):

Show me one step at a time, and I'll guide you through it.

OpenStudy (destinymasha):

4 x -3 - 5 + 2 . 4 x -3 = -12 x 2 = 24 24 - 5 = 19 + 2 = 21

OpenStudy (destinymasha):

wait.... i think i made a mistake

OpenStudy (mathstudent55):

F(x) = 2 |4x - 5| + 2 First, substitute -3 in for x: F(-3) = 2 | 4(-3) - 5 | + 2

OpenStudy (destinymasha):

4 x -3 = 12 , - 5 = 7 x 2 = 14, + 2 = 16?

OpenStudy (mathstudent55):

Now calculate the expression inside the absolute value sign: F(-3) = 2 | 4(-3) - 5 | + 2 F(-3) = 2 | -12 - 5 | + 2 F(-3) = 2 | -17 | + 2

OpenStudy (mathstudent55):

Now calculate the absolute value of -17: |-17| = 17, so F(-3) = 2 | -17 | + 2 F(-3) = 2(17) + 2 Now just multiply 17 by 2 and then add 2: F(-3) = 34 + 2 F(-3) = 36

OpenStudy (mathstudent55):

4 x (-3) = -12, not 12

OpenStudy (destinymasha):

ok, i wasn't calculating the absolute value. THX so much, now i know what i did wrong. [:

OpenStudy (destinymasha):

i thought i put - 12

OpenStudy (destinymasha):

guess not [:

OpenStudy (mathstudent55):

Then -12 - 5 = -17 then the absolute value of -17 is positive 17

OpenStudy (destinymasha):

yes, like 12 + 5

OpenStudy (destinymasha):

Thank you

OpenStudy (mathstudent55):

wlcm

OpenStudy (destinymasha):

Choose the correct graph for the following inequality. y ≤ 2x - 4

OpenStudy (mathstudent55):

Dashed lines mean simply < or >. You have <=, so it must be a solid line.

OpenStudy (mathstudent55):

Now you need to choose between the first and third graphs.

OpenStudy (mathstudent55):

Choose point (0, 0). Does point (0, 0) make the inequality true or false?